Compute the after - tax interest expense for a frim with interest on Excess Cash =$2000, Interest on Debt = $7000, and tax rate of 30%
Answer #1

Interest on debt is the interest expense which is $7,000. interest expense can be deducted from income before interest and tax to reduce tax expenses. So after-tax interest expense is interest*(1-tax rate) = $7,000*(1-0.30) = $7,000*0.70 = $4,900.

But we have received the interest on Excess Cash of $2,000 which will reduce interest expense. so, net after - tax interest expense is ($7,000 - $2,000)*(1-0.30) = $5,000*0.70 = $1,500.